# DirectorLLM for Human-Centric Video Generation

### v1

@@ 4.2 Training and Implementation Detail
Our training pipeline consists of four steps: (1) training a 6 depth residual vector-quantized VQ-VAE [38] human poses, effectively discretizing the human poses, (2) training the large language model Director on these pose tokens, (3) training a linear diffusion model for smoothing and interpolation, and (4) training a ControlNet [49] for the video generator.
As shown in Fig. 9 left part, to convert continuous human poses detected by OpenPose into discrete tokens compatible with our large language model, we train a temporally-aware vector-quantized variational autoencoder (VQ-VAE [38]). Our VQ-VAE is structured as a residual linear model with a depth of six, meaning it employs six codebooks, each with a size of 512. It takes in 20 pose key frames and convert them into 120 token ids. This configuration allows for high-fidelity discretization, capturing complex pose dynamics while efficiently encoding temporal information across frames. By transforming continuous pose vectors into discretized tokens, this VQVAE enables our large language model to effectively handle and predict pose sequences as structured data within a standard text generation pipeline.
Our large language model, fine-tuned from Llama 3 [10], is adapted to generate human poses tokens based on text prompts, with an extended vocabulary to accommodate newly added pose tokens. We show its structure in Fig. 9 middle part. The model is initialized from Llama 3 8B model checkpoint and all model parameters within its transformer stack are trained at this stage. The model is trained as a standard language generation model with PyTorch Lightning for 8 hours on 8 Nvidia A100 GPUs.
As shown in Fig. 9 right part, our linear interpolator integrates key frame poses and CLIP-encoded text prompts, producing smooth poses. The video ControlNet [49] is trained on dense pose frames where we randomly subsample 16 frames and its poses for training. We initialize it by duplicating the encoder structure and parameters of the UNet in VideoCrafter, as shown in Fig. 2. We keep the base UNet frozen during training and fine-tune only the pose ControlNet and zero conv layers. This approach allows our model to learn pose-specific control while maintaining the generative capabilities of the base video model.
@@ 4.3 Evaluation
Baselines and Evaluation setup Since we are the first work to train a LLM as director for human centric-video generation, we choose to compare with a few popular text to video generation models, including VideoCrafter [8], LaVie-2 [42], and CogVideo [17]. We craft a test dataset with 150 human-centric text prompts, focusing on cases with dynamic scenes and human poses.
Figure 6: Human evaluations show user preference towards our method across various aspects. Our model outperformed baselines in key metrics, particularly in Pose Alignment and Human Motion, demonstrating its superior capability in generating human-centric and text-aligned content.
4.3.1 Qualitative Evaluation
In our qualitative evaluation, we show outputs from each model for six video generation scenarios, as shown in Fig. 5. The baselines sometimes struggle with anatomical realism and temporal consistency, as indicated by several noticeable errors (highlighted in dashed colored boxes), including extra or broken limbs, and abrupt directional changes during movement. These issues reflect the baselines’ limitations in handling complex human dynamics, leading to unrealistic and visually jarring results. In contrast, our model avoids these visual artifacts, showing its improved performance in generating high-quality, human-centric videos. Due to space limits, we show more visual examples of generated pose and video frames from our model in Fig. 4 and in Appendix B.
Further more, our DirectorLLM generates 20 key pose frames which are then interpolated to 200 frames by pose interpolation model, sufficient for a video for as long as 7 seconds. We show examples of generate pose sequences and rendered RGB frames in Appendix C, demonstrating its capability of understand human dynamics and generating poses with high temporal consistency, dynamics and realism over extended durations.
4.3.2 Quantitative Evaluation
We run quantitative comparison on the entire test dataset and show the results as follows. In our experiments, we utilize VideoScore [15] to evaluate our text-to-video model. VideoScore is an automatic video quality assessment metric trained on VideoFeedback, a large-scale dataset containing human-provided scores for 37.6K synthesized videos from various generative models. Given its high correlation with human judgments, VideoScore serves as a reliable proxy for human evaluation, allowing us to comprehensively assess our model’s performance across these crucial aspects. We show the results in Tab. 1 for key evaluation dimensions: Visual Quality (VC), Temporal Consistency (TC), Dynamic Degree (DD), Text-to-Video Alignment (CLIPvid), and Factual Consistency (FC). Through quantitative evaluations, we show that our proposed method outputs realistic motion more diverse than baselines, especially in Text-to-Video Alignment (CLIPvid) score.
@@ 4.3.2 Quantitative Evaluation
4.3.3 Human Evaluation
Since there is no automatic video metric on the quality of human pose dynamic, we apply human evaluation specifically for this purpose. We compare the best results from each method and make a questionnaire for human evaluators. Specifically, we run blind A/B comparisons between videos generated from each model and ask questions on: (1) overall quality (Overall), (2) prompt alignment for the human pose (Pose Alignment), (3) prompt alignment for the video (Text Alignment), (4) generation quality for the dynamic human motion (Human Motion), and (5) generation quality for static human balance (Human Balance). For Text Alignment, annotators are instructed to consider all prompt specifications and the better one would be one that achieves all specifications in the prompt. Human Motion evaluates the realism of dynamic human movement within the generated video, focusing on natural and coherent motion patterns. Human Balance assesses the static realism of human figures, ensuring accurate proportions and anatomical balance, such as correct head-to-body ratios and plausible body alignment. For each pair of comparisons, we would randomize A/B and use 3 raters per comparison, and use the majority vote as the final decision. The human evaluation setup was inspired by the state-of-the-art video generation model [29].
Fig. 6 shows the results of human evaluation comparing our model to baselines (LaVie [42], CogVideo [17], and VideoCrafter [8]) across five categories. Our model consistently outperformed baselines in key metrics, particularly in Pose Alignment and Human Motion, demonstrating its superior capability in generating human-centric and text-aligned video content. This aligns with our core motivation to improve the realism and accuracy of human representations in generated videos by offloading high-level scene understanding and pose prediction to a dedicated language model. This enhances both the static and dynamic realism of human figure, fulfilling our objective to produce better human-centric videos.
@@ 4.4 Ablation and Analysis
Residual vector quantization. In this experiment, we ablate our design choice of residual pose VQ-VAE. The goal is to evaluate whether our 6 depth residual autoencoding benefits LLM’s predictive performance and video quality. As shown in Tab. 2, without residual autoencoding, the VQ-VAE couldn’t accurately reconstruct human pose, significantly degrading pose prediction accuracy and introducing inconsistencies in generated videos. Residual structure in our VQ-VAE, with a depth of 6, plays a crucial role in preserving pose fidelity, directly enhancing both LLM prediction accuracy and overall video quality.
Linear diffusion model. To assess the impact of our diffusion-based pose interpolation model, we replaced it with direct pose interpolation by linearly interpolating the 20 key frames to generate all 200 frames. These frames are then rendered to videos with our video generator. As shown in Tab. 2, the direct interpolation lacks authentic human dynamics and affects scores such as dynamic degree. Linear diffusion model outperformed basic interpolation in maintaining temporal consistency and generating smooth motion transitions, particularly in scenes with dynamic human movements.
Text condition in linear diffusion model. We also tested the effect of adding text-conditioned CLIP embeddings in the linear diffusion model by training it with and without the text conditions. As shown in Tab. 2, without text conditioning, the generated videos exhibited reduced alignment with the original text prompts, leading to inconsistencies in scene layout and subject interaction.
VC $\uparrow$ TC $\uparrow$ DD $\uparrow$ CLIPvid $\uparrow$ FC $\uparrow$
@@
Table 2: Comparison of different models across various metrics. Our ablation studies demonstrate the importance of each component in our model. Residual vector quantization improves pose reconstruction and LLM predictions, the linear diffusion model ensures temporal consistency and smooth motion, and text conditioning enhances alignment with prompts.
4.5 Limitations
While the DirectorLLM excels at predicting human pose dynamics, it does not consider interactions with non-human objects (e.g., vehicles, obstacles, or furniture) that could affect movement patterns. Without accounting for these elements, the model may produce human poses that appear unnatural or disconnected from the surrounding environment, as shown in Fig. 7. A more advanced model should take more scene details into consideration.
Figure 7: limitations. First row: The human in the video is running, but the background remains stationary. Second row: The human runs forward, but the background also moves forward.
@@ A. Implementation Details
The first sentence, known as the "Subject-Level Prompt", provides an overall video description with key information on human motions and poses. We use it as input to DirectorLLM transformer and Pose Linear Interpolator. We use full prompt as input to ControlNet Diffusion generator as it offers detailed descriptions of appearance and scene elements. We use OpenPose to extract per-frame human-poses for 25K videos. For videos with more than one human pose detected, we run a light weight human detection model InsightFace [12] to keep the one with the largest area and ignore all others.
@@
DirectorLLM. We leverage the advanced pre-existing knowledge from pretrained large languare model to aid in model convergence. Specifically, we initialize our DirectorLLM with Llama 3 -8B checkpoint[10] from HuggingFace. We extend its tokenizer and embeddings to an additonal 6 x 512 tokens, each of which representing a pose token from VQ-VAE codebooks. We attach 6 x 20 = 120 pose tokens from VQ-VAE to the "subject-level prompt", forming a sequence as long as 300 tokens. Training code is built with Pytorch-Lightining and the model is trained as a standard language generation model for 8 hours on 8 Nvidia A100 GPUs.
Pose Interpolation. To densify human motion from 1 fps to 30 fps, we construct an conditional diffusion model. It is built on top of the Audio2PhotoReal [25]. We follow the standard DDPM definition of diffusion, where we add noise to pose vectors during forward pass and train the network to remove it during as the denoising pass. It takes in noise and predict 200 pose vectors, each for one frame, conditoned on 20 key frame poses. To incorporate the text prompt information, we use CLIP text encoder and apply a cross attention layer. Timestep information is incorporated with a feature-wise linear modulation (FiLM) layer, similar to Audio2PhotoReal [25]. All parameters are optimized. Training is finished in 8 hours on one single A100 GPU.
Text-to-video ControlNet. We build our video renderer model on top of VideoCrafter2 [8]. A pose ControlNet [49] module is initialized by copying its UNet encoder structure and parameters. We convert poses for 16 frames from 32 dimensional vector to image, and feed it to ControlNet. Connection layers are initialzed with zero convolution for better converging. We lock the parameters of base UNet and optimize only our pose ControlNet. We apply large gradient accumulation steps (200) and clip grad norm for smoother training. Training is finished with AdamW optimizer and a learning rate of $1^{-5}$ on 32 A100 GPUs in 48 hours.
@@
We use FIFO [20] to bridge the gap between short-frame limitations and long-sequence requirements, which will be covered in section C.
@@ B. More Results from our model
We present additional visual results to demonstrate the effectiveness of our model in this section. Fig. 9 upper panel showcases the sequence of human poses predicted by the LLM Director, highlighting its ability to generate accurate and temporally coherent pose layouts based on text prompts. Fig. 8 pairs these pose layouts with their corresponding video renderings, illustrating how the poses guide the video generator to produce realistic, high-quality human motions and interactions. The alignment between the predicted poses and the rendered visuals demonstrates the strength of our multi-stage pipeline
@@ C. Extending Video Length
Our pose interpolation model generates 200 pose frames, while VideoCrafter with ControlNet is limited to generating 16 frames at a time. To extend the length of the generated video and render all 200 frames, we apply the FIFO-Diffusion [20] approach. This method enables infinite-length video generation by employing diagonal denoising, which processes consecutive frames in a queue. During this process, fully denoised frames at the head of the queue are dequeued, while new noise frames are enqueued at the tail for processing. To mitigate the training-inference gap caused by this iterative strategy, FIFO incorporates latent partitioning and lookahead denoising techniques [20], ensuring consistency across extended sequences. FIFO uses the exact same model checkpoint as our base model. Simply adding ControlNet modules to the FIFO code allows us to generate long videos with pose condition awareness, leveraging VideoCrafter’s capabilities to produce human-centric videos for as long as 200 frames. A sample is attached in Fig. 9 lower panel.

### v2

@@ 3.2 Training and Implementation Detail
Our training pipeline consists of four steps: (1) training a 6 depth residual vector-quantized VQ-VAE Van Den Oord et al. (2017) human poses, effectively discretizing the human poses, (2) training the large language model Director on these pose tokens, (3) training a linear diffusion model for smoothing and interpolation, and (4) training a ControlNet Zhang et al. (2023) for the video generator.
As shown in Fig. 8 left part, to convert continuous human poses detected by OpenPose into discrete tokens compatible with our large language model, we train a temporally-aware vector-quantized variational autoencoder (VQ-VAE Van Den Oord et al. (2017)). Our VQ-VAE is structured as a residual linear model with a depth of six, meaning it employs six codebooks, each with a size of 512. It takes in 20 pose key frames and convert them into 120 token ids. This configuration allows for high-fidelity discretization, capturing complex pose dynamics while efficiently encoding temporal information across frames. By transforming continuous pose vectors into discretized tokens, this VQVAE enables our large language model to effectively handle and predict pose sequences as structured data within a standard text generation pipeline.
Our large language model, fine-tuned from Llama 3 Dubey et al. (2024), is adapted to generate human poses tokens based on text prompts, with an extended vocabulary to accommodate newly added pose tokens. We show its structure in Fig. 8 middle part. The model is initialized from Llama 3 8B model checkpoint and all model parameters within its transformer stack are trained at this stage. The model is trained as a standard language generation model with PyTorch Lightning for 8 hours on 8 Nvidia A100 GPUs.
As shown in Fig. 8 right part, our linear interpolator integrates key frame poses and CLIP-encoded text prompts, producing smooth poses. The video ControlNet Zhang et al. (2023) is trained on dense pose frames where we randomly subsample 16 frames and its poses for training. We initialize it by duplicating the encoder structure and parameters of the UNet in VideoCrafter, as shown in Fig. 2. We keep the base UNet frozen during training and fine-tune only the pose ControlNet and zero conv layers. This approach allows our model to learn pose-specific control while maintaining the generative capabilities of the base video model.
@@ 3.3 Evaluation
Baselines and evaluation setup. Since we are the first work to train a LLM as director for human centric-video generation, we choose to compare with a few popular text to video generation models, including VideoCrafter Chen et al. (2023), LaVie-2 Wang et al. (2023b), and CogVideo Hong et al. (2022). We craft a test dataset with 150 human-centric text prompts, focusing on cases with dynamic scenes and human poses.
Qualitative evaluation. In our qualitative evaluation, we show outputs from each model for six video generation scenarios, as shown in Fig. 5. The baselines sometimes struggle with anatomical realism and temporal consistency, as indicated by several noticeable errors (highlighted in dashed colored boxes), including extra or broken limbs, and abrupt directional changes during movement. These issues reflect the baselines’ limitations in handling complex human dynamics, leading to unrealistic and visually jarring results. In contrast, our model avoids these visual artifacts, showing its improved performance in generating high-quality, human-centric videos. Due to space limits, we show more visual examples of generated pose and video frames from our model in Fig. 4. Further more, our DirectorLLM generates 20 key pose frames which are then interpolated to 200 frames by pose interpolation model, sufficient for a video for as long as 7 seconds. We show examples of generate pose sequences and rendered RGB frames in the supplementary, demonstrating its capability of understand human dynamics and generating poses with high temporal consistency, dynamics and realism over extended durations.
Quantitative evaluation.
Figure 6: Human evaluations show user preference towards our method.
We run a quantitative comparison on the entire test dataset and show the results as follows. In our experiments, we utilize VideoScore He et al. (2024b) to evaluate our text-to-video model. VideoScore is an automatic video quality assessment metric trained on VideoFeedback, a large-scale dataset containing human-provided scores for 37.6K synthesized videos from various generative models. Given its high correlation with human judgments, VideoScore serves as a reliable proxy for human evaluation, allowing us to comprehensively assess our model’s performance across these crucial aspects. We show the results in Fig. 1 for key evaluation dimensions: Visual Quality (VC), Temporal Consistency (TC), Dynamic Degree ($D$), Text-to-Video Alignment (CLIPvid), and Factual Consistency (FC). Through quantitative evaluations, we show that our proposed method outputs realistic motion more diverse than baselines, especially in Text-to-Video Alignment (CLIPvid) score.
@@
Human evaluation. Since there is no automatic video metric on the quality of human pose dynamic, we apply human evaluation specifically for this purpose. We compare the best results from each method and make a questionnaire for human evaluators. Specifically, we run blind A/B comparisons between videos generated from each model and ask questions on: (1) overall quality (Overall), (2) prompt alignment for the human pose (Pose Alignment), (3) prompt alignment for the video (Text Alignment), (4) generation quality for the dynamic human motion (Human Motion), and (5) generation quality for static human balance (Human Balance). For Text Alignment, annotators are instructed to consider all prompt specifications and the better one would be one that achieves all specifications in the prompt. Human Motion evaluates the realism of dynamic human movement within the generated video, focusing on natural and coherent motion patterns. Human Balance assesses the static realism of human figures, ensuring accurate proportions and anatomical balance, such as correct head-to-body ratios and plausible body alignment. For each pair of comparisons, we would randomize A/B and use 3 raters per comparison, and use the majority vote as the final decision. The human evaluation setup was inspired by the video generation model Polyak et al. (2024).
Fig. 6 shows the results of human evaluation comparing our model to baselines (LaVie Wang et al. (2023b), CogVideo Hong et al. (2022), and VideoCrafter Chen et al. (2023)) across five categories. Our model consistently outperformed baselines in key metrics, particularly in Pose Alignment and Human Motion, demonstrating its superior capability in generating human-centric and text-aligned video content. This aligns with our core motivation to improve the realism and accuracy of human representations in generated videos by offloading high-level scene understanding and pose prediction to a dedicated language model. This enhances both the static and dynamic realism of human figure, fulfilling our objective to produce better human-centric videos.
@@ 3.4 Ablation and Analysis
Residual vector quantization. In this experiment, we ablate our design choice of residual pose VQ-VAE. The goal is to evaluate whether our 6 depth residual autoencoding benefits LLM’s predictive performance and video quality. As shown in Fig. 2, without residual autoencoding, the VQ-VAE couldn’t accurately reconstruct human pose, significantly degrading pose prediction accuracy and introducing inconsistencies in generated videos. Residual structure in our VQ-VAE, with a depth of 6, plays a crucial role in preserving pose fidelity, directly enhancing both LLM prediction accuracy and overall video quality.
VC $\uparrow$ TC $\uparrow$ $D$ $\uparrow$ CLIPvid $\uparrow$ FC $\uparrow$
@@
Table 2: Comparison of different models across various metrics. Our ablation demonstrates the importance of each component. Residual vector quantization improves pose reconstruction and LLM predictions, the linear diffusion model ensures temporal consistency and smooth motion, and text conditioning enhances alignment with prompts.
Linear diffusion model. To assess the impact of our diffusion-based pose interpolation model, we replaced it with direct pose interpolation by linearly interpolating the 20 key frames to generate all 200 frames. These frames are then rendered to videos with our video generator. As shown in Table. 2, the direct interpolation lacks authentic human dynamics and affects scores such as dynamic degree. Linear diffusion model outperformed basic interpolation in maintaining temporal consistency and generating smooth motion transitions, particularly in scenes with dynamic human movements.
Text condition in linear diffusion model. We also tested the effect of adding text-conditioned CLIP embeddings in the linear diffusion model by training it with and without the text conditions. As shown in Table. 2, without text conditioning, the generated videos exhibited reduced alignment with the original text prompts, leading to inconsistencies in scene layout and subject interaction.
4 Discussions
As far as our best knowledge goes, we are the first work that have showcased the effectiveness of offloading human motions simulation to an LLM for human centric T2V generation. Our approach addresses the broader challenge of human video generation with dedicated motion reasoning. We believe this opens a way to motion-simulated video generators, where the motion and interaction of the main characters are driven by dedicated AIs.
@@ A. Implementation Details
The first sentence, known as the "Subject-Level Prompt", provides an overall video description with key information on human motions and poses. We use it as input to DirectorLLM transformer and Pose Linear Interpolator. We use full prompt as input to ControlNet Diffusion generator as it offers detailed descriptions of appearance and scene elements. We use OpenPose to extract per-frame human-poses for 25K videos. For videos with more than one human pose detected, we run a light weight human detection model InsightFace Guo and Deng (2024) to keep the one with the largest area and ignore all others.
@@
DirectorLLM. We leverage the advanced pre-existing knowledge from pretrained large languare model to aid in model convergence. Specifically, we initialize our DirectorLLM with Llama 3 -8B checkpointDubey et al. (2024) from HuggingFace. We extend its tokenizer and embeddings to an additonal 6 x 512 tokens, each of which representing a pose token from VQ-VAE codebooks. We attach 6 x 20 = 120 pose tokens from VQ-VAE to the "subject-level prompt", forming a sequence as long as 300 tokens. Training code is built with Pytorch-Lightining and the model is trained as a standard language generation model for 8 hours on 8 Nvidia A100 GPUs.
Pose Interpolation. To densify human motion from 1 fps to 30 fps, we construct an conditional diffusion model. It is built on top of the Audio2PhotoReal Ng et al. (2024). We follow the standard DDPM definition of diffusion, where we add noise to pose vectors during forward pass and train the network to remove it during as the denoising pass. It takes in noise and predict 200 pose vectors, each for one frame, conditoned on 20 key frame poses. To incorporate the text prompt information, we use CLIP text encoder and apply a cross attention layer. Timestep information is incorporated with a feature-wise linear modulation (FiLM) layer, similar to Audio2PhotoReal Ng et al. (2024). All parameters are optimized. Training is finished in 8 hours on one single A100 GPU.
Text-to-video ControlNet. We build our video renderer model on top of VideoCrafter2 Chen et al. (2023). A pose ControlNet Zhang et al. (2023) module is initialized by copying its UNet encoder structure and parameters. We convert poses for 16 frames from 32 dimensional vector to image, and feed it to ControlNet. Connection layers are initialzed with zero convolution for better converging. We lock the parameters of base UNet and optimize only our pose ControlNet. We apply large gradient accumulation steps (200) and clip grad norm for smoother training. Training is finished with AdamW optimizer and a learning rate of $1^{-5}$ on 32 A100 GPUs in 48 hours.
@@
We use FIFO Kim et al. (2024) to bridge the gap between short-frame limitations and long-sequence requirements, which will be covered in section C.
@@ B. More Results from Our Model
We present additional visual results to demonstrate the effectiveness of our model in this section. figure˜8 upper panel showcases the sequence of human poses predicted by the LLM Director, highlighting its ability to generate accurate and temporally coherent pose layouts based on text prompts. figure˜7 pairs these pose layouts with their corresponding video renderings, illustrating how the poses guide the video generator to produce realistic, high-quality human motions and interactions. The alignment between the predicted poses and the rendered visuals demonstrates the strength of our multi-stage pipeline
@@ C. Extending Video Length
Our pose interpolation model generates 200 pose frames, while VideoCrafter with ControlNet is limited to generating 16 frames at a time. To extend the length of the generated video and render all 200 frames, we apply the FIFO-Diffusion Kim et al. (2024) approach. This method enables infinite-length video generation by employing diagonal denoising, which processes consecutive frames in a queue. During this process, fully denoised frames at the head of the queue are dequeued, while new noise frames are enqueued at the tail for processing. To mitigate the training-inference gap caused by this iterative strategy, FIFO incorporates latent partitioning and lookahead denoising techniques Kim et al. (2024), ensuring consistency across extended sequences. FIFO uses the exact same model checkpoint as our base model. Simply adding ControlNet modules to the FIFO code allows us to generate long videos with pose condition awareness, leveraging VideoCrafter’s capabilities to produce human-centric videos for as long as 200 frames. A sample is attached in figure˜8 lower panel.